A lady parks her car on the 6th floor and took a lift up 17 floors to the finance department. She went down 9 floors to the tax department. On which floor is the lady on
1
Expert's answer
2019-12-23T09:10:42-0500

Initial position:


"x(0)=6"

The finance department is at


"x(1)=6+17=23"

Final position:


"x(2)=23-9=14"

The lady is on 14th floor.